# Light deflection in axially symmetric stationary spacetimes filled with a moving medium

### v1

@@ II Previous studies of light propagation in axially symmetric spacetimes in presence of cold plasma
Crisnejo et al. [49] applied the Gauss-Bonnet theorem to calculate the deflection angles in stationary axially symmetric spacetimes in the presence of a cold plasma. In particular, they derived the deflection angle for light rays in the equatorial plane of the Kerr metric in the weak deflection case, including higher-order terms. To study the related works of this group with other interesting results, see also [46, 47, 48, 63, 96]. Bezděková and Bičák [104] studied the deflection angle in the equatorial plane around an axially symmetric stationary object emersed in a general static dispersive medium characterized solely by its refractive index. Within this setting, they derived an integral expression for the deflection angle and applied it to the case of a cold plasma in the Hartle-Thorne metric. Additionally, they performed a weak deflection analysis and compared results between the Hartle-Thorne and Kerr spacetimes.
@@ III Synge’s formalism and equations of motion with general refractive index
$\mathcal{H}(x,p)=\frac{1}{2}\left\{g^{ik}p_{i}p_{k}-\left[n^{2}(x\omega(x,p))-1\right]\left(p_{j}V^{j}\right)^{2}\right\}\,$, (1)
@@
The gravitational field is given by the spacetime metric $g^{ik}$, while the medium is characterized by its refractive index $n=n(x\omega)$ and its 4-velocity $V^{j}=V^{j}(x)$ satisfying
@@
See also the discussion of this metric in the Appendix of [93]. The spherically symmetric case was considered in our previous paper [105], and can be reconstructed by setting $P(r,\vartheta)\to 0$, $A(r,\vartheta)\to A(r)$, $B(r,\vartheta)\to B(r)$ and $C(r,\vartheta)\to D(r)\sin^{2}\vartheta$, $D(r,\vartheta)\to D(r)$.
@@ IV.2 Deflection Angle for a Radially Moving Medium
The deflection angle calculated in this manner is always positive, regardless of the direction of $\varphi$-motion. One should use the “plus” sign in front of $h(R)$ in (47) for motion with increasing $\varphi$ (i.e., $\dot{\varphi}>0$), and the “minus” sign when $\varphi$ is decreasing ($\dot{\varphi}<0$). For quick reference, we recall that the coefficients $A(r)$, $B(r)$, $C(r)$ and $P(r)$ are the metric functions (cf. (9)); the expression $\mathcal{A}_{r}(r)$ is given by (23); and the function $h(r)$ is defined by (39). The deflection angle depends on two additional parameters: the closest approach distance $R$ and the photon frequency at infinity $\omega_{0}$; the latter being embedded in function $h(r)$.
@@ V.1 The Synge Hamiltonian to linear order in the medium velocity on arbitrary spacetime
Note that only one combination of signs was left here, in accordance with the discussion below Eq. (21). We also recall that the metric component $g_{00}$ is negative, corresponding to the signature convention (,+,+,+)(-,+,+,+).
@@
    $=n(x,\bar{\omega})+\left(\frac{\partial n}{\partial\omega}(x\bar{\omega})\right)\left[\left(\frac{g_{0\alpha}p_{0}}{g_{00}}-p_{\alpha}\right)V^{\alpha}\right]$
@@
    $+\frac{2p_{0}}{g_{00}^{2}}\left[p_{0}\ n(x,\bar{\omega})\left(\tfrac{\partial n}{\partial\omega}(x\bar{\omega})\right)-\sqrt{-g_{00}}\,(n(x,\bar{\omega})^{2}-1)\right]$
@@ Acknowledgments
CP acknowledges support by the excellence cluster QuantumFrontiers of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967 and was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project Number 420243324.
BB thanks to Prof. Pavel Krtouš and the Institute of Theoretical Physics at the Charles University in Prague for supporting her research during autumn 2024, when the core results of this paper were derived.

### v2

@@ II Previous studies of light propagation in axially symmetric spacetimes in presence of cold plasma
Crisnejo et al. [49] applied the Gauss-Bonnet theorem to calculate the deflection angles in stationary axially symmetric spacetimes in the presence of a cold plasma. In particular, they derived the deflection angle for light rays in the equatorial plane of the Kerr metric in the weak deflection case, including higher-order terms. To study the related works of this group with other interesting results, see also [46, 47, 48, 63, 96]. Bezděková and Bičák [104] studied the deflection angle in the equatorial plane around an axially symmetric stationary object immersed in a general static dispersive medium characterized solely by its refractive index. Within this setting, they derived an integral expression for the deflection angle and applied it to the case of a cold plasma in the Hartle-Thorne metric. Additionally, they performed a weak deflection analysis and compared results between the Hartle-Thorne and Kerr spacetimes.
@@ III Synge’s formalism and equations of motion with general refractive index
$\mathcal{H}(x,p)=\frac{1}{2}\left\{g^{ik}p_{i}p_{k}-\left[n^{2}(x,\omega(x,p))-1\right]\left(p_{j}V^{j}\right)^{2}\right\}\,$, (1)
@@
The gravitational field is given by the spacetime metric $g^{ik}$, while the medium is characterized by its refractive index $n=n(x,\omega)$ and its 4-velocity $V^{j}=V^{j}(x)$ satisfying
@@
See also the discussion of this metric in the Appendix of [93]. The spherically symmetric case was considered in our previous paper [105], and can be reconstructed by setting $P(r,\vartheta)\to 0$, $A(r,\vartheta)\to A(r)$, $B(r,\vartheta)\to B(r)$, and $C(r,\vartheta)\to D(r)\sin^{2}\vartheta$, $D(r,\vartheta)\to D(r)$.
@@ IV.2 Deflection Angle for a Radially Moving Medium
The deflection angle calculated in this manner is always positive, regardless of the direction of $\varphi$-motion. One should use the “plus” sign in front of $h(R)$ in (47) for motion with increasing $\varphi$ (i.e., $\dot{\varphi}>0$), and the minus sign when $\varphi$ is decreasing ($\dot{\varphi}<0$). For quick reference, we recall that the coefficients $A(r)$, $B(r)$, $C(r)$ and $P(r)$ are the metric functions (cf. (9)); the expression $\mathcal{A}_{r}(r)$ is given by (23); and the function $h(r)$ is defined by (39). The deflection angle depends on two additional parameters: the closest approach distance $R$ and the photon frequency at infinity $\omega_{0}$; the latter being embedded in function $h(r)$.
@@ V.1 The Synge Hamiltonian to linear order in the medium velocity on arbitrary spacetime
Note that only one combination of signs was left here, in accordance with the discussion below Eq. (21). We also recall that the metric component $g_{00}$ is negative, corresponding to the signature convention ${,+,+,+}\{$-,+,+,+\}.
@@
    $=n(x,\bar{\omega})+\left(\frac{\partial n}{\partial\omega}(x,\bar{\omega})\right)\left[\left(\frac{g_{0\alpha}p_{0}}{g_{00}}-p_{\alpha}\right)V^{\alpha}\right]$
@@
    $+\frac{2p_{0}}{g_{00}^{2}}\left[p_{0}\ n(x,\bar{\omega})\left(\tfrac{\partial n}{\partial\omega}(x,\bar{\omega})\right)-\sqrt{-g_{00}}\,(n(x,\bar{\omega})^{2}-1)\right]$
@@ Acknowledgments
C.P. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967 and was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project Number 420243324.
B.B. thanks Professor Pavel Krtouš and the Institute of Theoretical Physics at the Charles University in Prague for supporting her research during autumn 2024, when the core results of this paper were derived.
O.Y.T. thanks the Excellence Cluster QuantumFrontiers of the German Research Foundation for supporting his visit to ZARM, University of Bremen, and Eva Hackmann and her group for their kind hospitality.